# Primus: A Pioneering Collection of Open-Source Datasets for Cybersecurity LLM Training

### v1

@@ Abstract
Large Language Models (LLMs) have shown remarkable advancements in specialized fields such as finance, law, and medicine. However, in cybersecurity, we have noticed a lack of open-source datasets, with a particular lack of high-quality cybersecurity pretraining corpora, even though much research indicates that LLMs acquire their knowledge during pretraining. To address this, we present a comprehensive suite of datasets covering all major training stages, including pretraining, instruction fine-tuning, and reasoning distillation with cybersecurity-specific self-reflection data. Extensive ablation studies demonstrate their effectiveness on public cybersecurity benchmarks. In particular, continual pre-training on our dataset yields a 15.88% improvement in the aggregate score, while reasoning distillation leads to a 10% gain in security certification (CISSP). We will release all datasets and trained cybersecurity LLMs under the ODC-BY and MIT licenses to encourage further research in the community. [^1] For access to all datasets and model weights, please refer to this link.
@@ 1 Introduction
Common training methods for LLMs include pre-training (PT) Radford (2018), supervised fine-tuning (SFT) Zhang et al. (2023), and reinforcement learning (RL) Wang et al. (2024b). Recent studies suggest LLMs acquire knowledge primarily during PT, and continual pre-training (CPT) Wu et al. (2024), which further trains pre-trained models on large amounts of unlabeled domain-specific text, can enhance their grasp of domain knowledge. In contrast, SFT may introduce hallucinations as new knowledge is learned Gekhman et al. (2024). More recently, collecting reflection data from reasoning models for distillation has also become a trend Huang et al. (2024). Typically, obtaining a domain-specific LLM may require applying multiple training methods, as in our pipeline (Fig.1).
@@
The cybersecurity field has yet to fully benefit from this transformative technology, which requires domain expertise due to its broad and complex nature. Our statistics on cybersecurity LLM survey papers Zhang et al. (2024a); Xu et al. (2024) indicate that most existing research focuses on SFT to align model outputs, while PT or CPT is largely performed on non-natural language data such as assembly code Jiang et al. (2023); Wang et al. (2024a); Sun et al. (2023), as shown in Fig.2. Clearly, these approaches have limited effectiveness in improving the general cybersecurity knowledge of LLMs. On the other hand, models pre-trained on cybersecurity knowledge Park and You (2023); Ranade et al. (2021); Jackaduma (2021); Aghaei et al. (2022) are limited to small ones like BERT Devlin et al. (2019), and none of them have released datasets. To the best of our knowledge, LLMs pre-trained on cybersecurity knowledge or distilled on reasoning data from cybersecurity tasks remain unexplored.
@@
• A Collection of Cybersecurity Datasets. We create a series of carefully curated datasets covering multiple stages of LLM training, including pre-training (Primus-Pretraining), instruction fine-tuning (Primus-Instruct), and reasoning fine-tuning (Primus-Reasoning), as shown in Fig.1. Extensive ablation studies and evaluations on cybersecurity benchmarks show that these datasets can effectively improve cybersecurity capabilities. All datasets will be released under a ODC-BY license to encourage further research in the community.
• A Family of Cybersecurity LLMs. We present a family of cybersecurity LLMs designed to tackle domain-specific challenges, including Llama-Primus-Base, a model continually pre-trained with cybersecurity knowledge text based on Llama-3.1-8B-Instruct, achieving a 15.88% improvement on aggregated cybersecurity benchmarks; Llama-Primus-Merged, an instruction-tuned variant merged with Llama-3.1-8B-Instruct, which retains instruction-following capability while significantly improving cybersecurity performance; and Llama-Primus-Reasoning, which is distilled from reasoning steps with reflection generated by a larger reasoning LLM on cybersecurity tasks, providing it long-thought capabilities and yielding a 10% gain on security certification. Likewise, all models will be released under an MIT license.
@@ 2.2.1 Composition
We specifically collect cybersecurity-related text from diverse sources, including Blogs, News, Books, Websites, Wikipedia, and MITRE, guided by prior pretraining work Aghaei et al. (2022). For Blogs and News, we select content from government agencies, standards bodies, cybersecurity companies, media, and forums. Meanwhile, Books cover a wide range of cybersecurity topics, and we exclude covers, tables of contents, and appendices while treating each extracted page as a separate sample. We also collect Webpages from well-known cybersecurity companies, which may include product descriptions, company profiles, FAQs, and API documentation. In addition, Wikipedia does not provide a predefined cybersecurity subset, so we perform a custom filtering process. Each Wikipedia article is associated with one or more category tags, which can be further expanded into subcategory tags. Starting from the root category "Computer Security", we recursively traverse its subcategories, using GPT-4o to determine whether a category is cybersecurity-related [^3] The prompt is provided in the Appx.A (Fig.6). This process yields 375 relevant categories, from which we extract corresponding Wikipedia articles. For MITRE, we leverage obsidian-mitre-attack [^4] https://github.com/vincenzocaputo/obsidian-mitre-attack, which converts STIX data from the official repository into readable Markdown.
@@ 2.2.2 Preprocessing Pipeline
We use perplexity from a language model trained on English Wikipedia as a quality score. Specifically, we use a 5-gram KenLM language model Heafield (2011) due to its efficiency in processing large amounts of data. With this setup, we manually inspect and determine an appropriate perplexity threshold for each source, and remove texts whose perplexity exceeds the threshold.
@@
Deduplication has been correlated with improvements in model performance Lee et al. (2022). We adopt FineWeb’s deduplication strategy, using a fuzzy hash-based approach with MinHash, which scales efficiently across many CPU nodes and allows tuning of similarity thresholds by adjusting the number of hashes per bucket. Specifically, we extract the 5 grams of each document and compute MinHashes using 112 hash functions, divided into 14 buckets of 8 hashes each to target documents that are at least 75% similar. Documents that share the same 8 MinHashes in any bucket are considered duplicates.
@@ 2.2.3 Augmentation
We find that some web-scraped data contains valuable information but suffers from poor readability due to irregular formatting, such as inconsistent line breaks. To address this, we adopt a rewriting approach inspired by Cosmopedia [^5] https://github.com/huggingface/cosmopedia, a reproduction of the high-quality synthetic dataset used in phi-1.5 Li et al. (2023b). Specifically, we prompt an LLM to rewrite the given text into a specific style, including blog posts, textbooks, and Q&A formats [^6] The prompt is provided in the Appx.A (Fig.7). To increase diversity, the rewriting LLM is randomly selected from GPT-4o, Llama-3.1-405B-Instruct, DBRX Mosaic (2024), and Claude 3.5 Sonnet Anthropic (2024).
@@ 2.3.1 Cybersecurity Classifier
We then use the classifier to score all FineWeb texts on a scale from 0 to 1, where higher scores indicate greater cybersecurity relevance. The distribution in Fig.3 shows that lower scores correspond to a significant increase in text volume. To determine an appropriate threshold for filtering, we first verify that whether texts with higher scores are truly cybersecurity-related. To do this, we leverage GPT-4o for accurate evaluation by dividing the scores into multiple bins, with dynamically adjusted bin sizes—smaller bins for lower scores—to account for the increased volume of data in lower score ranges. We randomly sample 50 texts from each bin and prompt GPT-4o [^7] The prompt is provided in the Appx.A (Fig.8) for classification. As shown in Fig.4, relevant text proportions remain above 60% at higher scores, but drop below 50% when scores fall below 0.003. Although incorporating some general text can help mitigate catastrophic forgetting Sun et al. (2019), we prioritize maintaining a majority of cybersecurity content. Therefore, we set the final threshold at 0.003, which corresponds to 15.3B of FineWeb data.
@@ 2.4 Primus-Instruct
After pre-training, we use Primus-Instruct for instruction fine-tuning to restore the instruction-following capability of the model. To achieve this, we design several hundred cybersecurity tasks covering common business scenarios, including explaining detected alerts, answering questions about retrieved security documents, analyzing executed suspicious commands, generating query languages for retrieving security events, and providing security recommendations and risk assessments for Terraform configurations. Each example is answered by GPT-4o, and we further use Claude 3.5 Sonnet as a judge [^8] The judge prompt is provided in the Appx.A (Fig.9) to discard samples with insufficiently helpful answers. In addition, we include several hundred multi-turn conversations on general topics generated by GPT-4o. As a result, these form Primus-Instruct, with statistics in Tab.3.
@@ 2.5 Primus-Reasoning
With the release of OpenAI’s reasoning model o1, an increasing number of studies have attempted to replicate its reasoning capabilities. One widely recognized approach is distillation, where reasoning samples with self-reflection from existing reasoning models are used to guide models in acquiring long-thought capabilities Huang et al. (2024); Liu et al. (2024). To this end, we select the following cybersecurity reasoning tasks from CTI-Bench Alam et al. (2024) and prompt o1-preview one to two times per question to generate solutions with reasoning steps and reflection [^9] The prompt is provided in the Appx.A (Fig.10), applying rejection sampling to retain only the correctly answered samples. The dataset statistics are shown in Tab.4.
Dataset Total Accepted Avg. Tokens
CTI-MCQ 1000 806 691.67
CTI-RCM 1000 728 761.10
CTI-RCM-2021 1000 635 766.47
CTI-VSP 1000 231 1155.83
CTI-ATE 60 2 1313.50

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable advancements in specialized fields such as finance, law, and medicine. However, in cybersecurity, we have noticed a lack of open-source datasets, with a particular lack of high-quality cybersecurity pretraining corpora, even though much research indicates that LLMs acquire their knowledge during pretraining. To address this, we present a comprehensive suite of datasets covering all major training stages, including pretraining, instruction fine-tuning, and reasoning distillation with cybersecurity-specific self-reflection data. Extensive ablation studies demonstrate their effectiveness on public cybersecurity benchmarks. In particular, continued pre-training on our dataset yields a 15.9% improvement in the aggregate score, while reasoning distillation leads to a 15.8% gain in security certification (CISSP). We will release all datasets and trained cybersecurity LLMs under the ODC-BY and MIT licenses to encourage further research in the community. [^1] For access to all datasets and model weights, please refer to this link.
@@ 1 Introduction
Common training methods for LLMs include pre-training (PT) Radford (2018), supervised fine-tuning (SFT) Zhang et al. (2023b), and reinforcement learning (RL) Wang et al. (2024b). Recent studies suggest LLMs acquire knowledge primarily during PT, and continued pre-training (CPT) Gururangan et al. (2020), which further trains pre-trained models on large amounts of domain-specific text, can enhance their grasp of domain knowledge. In contrast, SFT may introduce hallucinations as new knowledge is learned Gekhman et al. (2024). More recently, collecting reflection data from reasoning models for distillation has also become a trend Huang et al. (2024). Typically, obtaining a domain-specific LLM may require applying multiple training methods, as in our pipeline (Fig.1).
@@
The cybersecurity field has yet to fully benefit from this transformative technology, which requires domain expertise due to its broad and complex nature. Our statistics on cybersecurity LLM survey papers Zhang et al. (2024a); Xu et al. (2024a) indicate that most existing research focuses on SFT to align model outputs, while PT or CPT is largely performed on non-natural language data such as assembly code Jiang et al. (2023); Wang et al. (2024a); Sun et al. (2023), as shown in Fig.2. Clearly, these approaches have limited effectiveness in improving the general cybersecurity knowledge of LLMs. On the other hand, models pre-trained on cybersecurity knowledge Park and You (2023); Ranade et al. (2021); Jackaduma (2021); Aghaei et al. (2022) are limited to small ones like BERT Devlin et al. (2019), and none of them have released datasets. To the best of our knowledge, LLMs pre-trained on cybersecurity knowledge or distilled on reasoning data from cybersecurity tasks remain unexplored.
@@
• A Collection of Cybersecurity Datasets. We create a series of carefully curated datasets covering multiple stages of LLM training, including pre-training (Primus-Pretraining), instruction fine-tuning (Primus-Instruct), and reasoning fine-tuning (Primus-Reasoning), as shown in Fig.1. Extensive ablation studies and evaluations on cybersecurity benchmarks show that these datasets can effectively improve cybersecurity capabilities. Alldatasets will be released under the ODC-BY license to encourage further research in the community.
• A Family of Cybersecurity LLMs. We present afamily of cybersecurity LLMs designed to tackle domain-specific challenges, including Llama-Primus-Base, a model further pre-trained with cybersecurity knowledge based on Llama-3.1-8B-Instruct, achieving a 15.9% improvement on aggregated cybersecurity benchmarks; Llama-Primus-Merged, an instruction-tuned variant merged with Llama-3.1-8B-Instruct, which retains general instruction-following capability while significantly improving cybersecurity performance; and Llama-Primus-Reasoning, which is distilled from reasoning steps with reflection generated by a larger reasoning LLM on cybersecurity tasks, providing it long-thought capabilities and yielding a 15.8% gain on security certification. Likewise, all models will be released under the MIT license.
@@ 2.2.1 Composition
We specifically collect cybersecurity-related text from diverse sources, including Blogs, News, Books, Websites, Wikipedia, and MITRE, guided by prior pretraining work Aghaei et al. (2022). For Blogs and News, we select content from government agencies, standards bodies, cybersecurity companies, media, and forums. Meanwhile, Books cover a wide range of cybersecurity topics, and we exclude covers, tables of contents, and appendices while treating each extracted page as a separate sample. We also collect Webpages from well-known cybersecurity companies, which may include product descriptions, company profiles, FAQs, and API documentation. In addition, Wikipedia does not provide a predefined cybersecurity subset, so we perform a custom filtering process. Each Wikipedia article is associated with one or more category tags, which can be further expanded into subcategory tags. Starting from the root category "Computer Security", we recursively traverse its subcategories, using GPT-4o to determine whether a category is cybersecurity-related [^3] The prompt is provided in the Appx.E (Fig.8). This process yields 375 relevant categories, from which we extract corresponding Wikipedia articles. For MITRE, we leverage obsidian-mitre-attack [^4] https://github.com/vincenzocaputo/obsidian-mitre-attack, which converts STIX data from the official repository into readable Markdown.
@@ 2.2.2 Preprocessing Pipeline
We use perplexity from a language model trained on English Wikipedia as a quality score. Specifically, we use a 5-gram KenLM language model Heafield (2011) due to its efficiency in processing large amounts of data. With this setup, we manually set an appropriate perplexity threshold for each source, and remove texts whose perplexity exceeds the threshold.
@@
Deduplication has been correlated with improvements in model performance Lee et al. (2022). We adopt FineWeb’s deduplication strategy, using a fuzzy hash-based approach with MinHash. Specifically, we extract 5-grams from each document and compute MinHashes using 112 hash functions, split into 14 buckets of 8 hashes each to target documents at least 75% similar. Documents sharing the same 8 MinHashes in any bucket are considered duplicates.
@@ 2.2.3 Augmentation
We find that some web-scraped data contains valuable information but suffers from poor readability due to irregular formatting, such as inconsistent line breaks. To address this, we adopt a rewriting approach inspired by Cosmopedia [^5] https://github.com/huggingface/cosmopedia, a reproduction of the high-quality synthetic dataset used in phi-1.5 Li et al. (2023b). Specifically, we prompt an LLM to rewrite the given text into a specific style, including blog posts, textbooks, and Q&A formats [^6] The prompt is provided in the Appx.E (Fig.9). To increase diversity, the rewriting LLM is randomly selected from GPT-4o, Llama-3.1-405B-Instruct, DBRX Mosaic (2024), and Claude 3.5 Sonnet Anthropic (2024).
@@ 2.3.1 Cybersecurity Classifier
We then use the classifier to score all FineWeb texts on a scale from 0 to 1, where higher scores indicate greater cybersecurity relevance. The distribution in Fig.3 shows that lower scores correspond to a significant increase in text volume. To determine an appropriate threshold for filtering, we first verify that whether texts with higher scores are truly cybersecurity-related. To do this, we leverage GPT-4o for accurate evaluation by dividing the scores into multiple bins, with dynamically adjusted bin sizes—smaller bins for lower scores—to account for the increased volume of data in lower score ranges. We randomly sample 50 texts from each bin and prompt GPT-4o [^7] The prompt is provided in the Appx.E (Fig.10) for classification. As shown in Fig.4, relevant text proportions remain above 60% at higher scores, but drop below 50% when scores fall below 0.003. Although incorporating some general text can help mitigate catastrophic forgetting Sun et al. (2019), we prioritize maintaining a majority of cybersecurity content. Therefore, we set the final threshold at 0.003, which corresponds to 15.3B of FineWeb data.
@@ 2.4 Primus-Instruct
After pre-training, we use Primus-Instruct for instruction fine-tuning to restore the instruction-following capability of the model. To achieve this, we design several hundred cybersecurity tasks covering common business scenarios, including explaining detected alerts, answering questions about retrieved security documents, analyzing executed suspicious commands, generating query languages for retrieving security events, and providing security recommendations and risk assessments for Terraform configurations. Each example is answered by GPT-4o, and we further use Claude 3.5 Sonnet as a judge [^8] The judge prompt is provided in the Appx.E (Fig.11) to discard samples with insufficiently helpful answers. In addition, we include several hundred multi-turn conversations on general topics generated by GPT-4o. As a result, these form Primus-Instruct, with statistics in Tab.3.
@@ 2.5 Primus-Reasoning
With the release of OpenAI’s reasoning model o1, an increasing number of studies have attempted to replicate its reasoning capabilities. One widely recognized approach is distillation, where reasoning samples with self-reflection from existing reasoning models are used to guide models in acquiring long-thought capabilities Huang et al. (2024); Liu et al. (2024). To this end, we select cybersecurity reasoning tasks from CTI-Bench [^9] A brief introduction to CTI-Bench is provided in Appx.C Alam et al. (2024) and prompt o1-preview one to two times per question to generate solutions with reasoning steps and reflection [^10] The prompt is provided in the Appx.E (Fig.12), applying rejection sampling to retain only the correctly answered samples. We also include DeepSeek-R1, obtained by directly querying its open-source model to access reasoning steps. The dataset statistics are shown in Tab.4.
Dataset Samples Accepted Avg. Tokens
(o1-preview / DeepSeek-R1)
CTI-MCQ 1000 806 / 768 692 / 672
CTI-RCM 1000 728 / 721 761 / 530
CTI-RCM-2021 1000 635 / 683 766 / 543
CTI-VSP 1000 231 / 312 1156 / 1395
CTI-ATE 60 2 / 5 1314 / 1731